OpenStudy (anonymous):

The members of the Student Catholic Action decided to raise $ 4000 to help the local charity clinic. The amount to be raised was divided equally among the members. Due to an increase of 20 members, the amount to be raised by each member went down by $10. How many members were originally in the club?

OpenStudy (anonymous):

amount per member, if you call the original number of members say \(x\) then the amount per member is \(\frac{4000}{x}\)

OpenStudy (anonymous):

if there are 20 more members then the amount per member is \(\frac{4000}{x+20}\)

OpenStudy (anonymous):

you are told that the second one is ten smaller than the first solve \[\frac{4000}{x}-10=\frac{4000}{x+20}\]

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

solve for \(x\)!

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

don't get how i got the equation? or how to solve for \(x\) ?

OpenStudy (anonymous):

both huhu sorry

OpenStudy (anonymous):

i wrote my best explanation above not sure how else to put it

OpenStudy (anonymous):

how to solve for x then? :(